A bag contains 36 marbles, some of which are red and the remainder of which are blue. Twice the number of red marbles is six less than the number of blue marbles. Find the number of marbles of each color. (Answer: 10 red marbles; 26 blue marbles). Just need the work to be shown. Thanks.
Answer by stanbon(75887)

A bag contains 36 marbles, some of which are red and the remainder of which are blue. Twice the number of red marbles is six less than the number of blue marbles. Find the number of marbles of each color.
---------------
Equations:
r + b = 36 marbles
2r = b-6
----
Solve for "b" and substitute:
b = 2r+6
---
r + 2r+6 = 36
3r = 30
r = 10 (# of red marbles)
----
b = 36-10 = 26 (# of blue marbles)
====================================
Cheers,
Stan H.